A recently married couple would like to start a family. They are both heterozygous for albinism, a recessive genetic disorder. They would like to have three children.

A. What is the probability that their first child would be albino? A. _____________________________

B. What is the probability that their first and second and third child would be albino? B. _____________________________

C. What is the chance that they will have 3 girls that are all albino? C. _____________________________

D. What is the chance that they will have 3 boys that are all albino? D. ________________
Answer #1

Suppose both the parents have Aa genotype.

Aa x Aa

A a
A AA Aa
a Aa aa

A) chance of the 1st child to be born with albino is 1/4.

B) The probability for each child being born is the same; therefore 1st, 2nd, and 3rd child to have albino is 1/4*1/4*1/4= 1/64.

C) The probability of having a girl = ½. The probability of having a girl with albino is ¼ * ½ = 1/8. The probability of having three girls with albino = 1/8 * 1/8 * 1/8= 1/512.

D) The probability of having a boy = ½. The probability of having a boy with albino is ¼ * ½ = 1/8. The probability of having three boys with albino = 1/8 * 1/8 * 1/8= 1/512.